OpenStudy (anonymous):

how do you find the domain of sqrt(x+4)-2?

OpenStudy (agreene):

Well, lets think to the left first. When x=-4 we have \[\sqrt{-4+4}-2=\sqrt 0 -2 = -2\] but when we go to x=-5 \[\sqrt{-5+4}-2=\sqrt{-1}-2=-2+i\] So, it isnt defined by the reals starting at x=-5, but actually thats true at x=-4.00001 Now we know, x=-4 is the furthest left we can go. Getting bigger, we dont have any threat of going undefined \[\sqrt{10^{100}-2}-2\] exists, so our domain must be \[-4\le x\le\infty\]

OpenStudy (anonymous):

I see now, I was missing an easy piece. Thank you!

OpenStudy (agreene):

no problem.